For the diprotic weak acid H2A, Ka1-2.8 x10 and Ka28.9 x109 What is the pH of a 0.0450 M solution of H2A? What are the equilibrium concentrations of H2A and A in this solution? Number Number Number 2-Help please
Answer #1

H2A <------> HA- + H+;..................Ka1 = 2.8x10-6
HA- <------> A2- + H+;...................Ka2 = 8.9x10-9

For the concentrations of H+ and HA- you consider only the first reaction since the HA- product will only slightly dissociate (Ka2 is very small).

If x moles of H2A dissociate, then x moles of HA- and H+ are produced, and the concentration of H2A is reduced by x:

[H2A] = 0.0450 - x
[H+] = x
[HA-] = x

Ka1 = [H+]x[HA-] / [H2A] = x² / (0.0450 - x)

2.8x10-6 x (0.0450 - x) = x²
x² + 2.8 x10-6x – (0.0450 x 2.8x10-6) = 0
x = 0.0017

[H+] = x
pH = -log(x) = 2.8
[H2A] = 0.045 - x = 0.043 M

For the second reaction, Ka2 = [A2-]x[H+] / [HA-] = ([A2-] x) / x;

   [A2-] = Ka2
   Therefore    [A2-]= 8.9*10-9 M